A particle is projected with a velocity of 19.6m/s at an angle of 30° with the horizontal. After what time will it move in a direction perpendicular to its original direction of motion?

Grade:11

1 Answers

Khimraj
5 years ago
initial velocity =
v1 = 19.6cos30\hat{i} +19.6sin30\hat{j} = 9.8\sqrt{3}\hat{i}  + 9.8\hat{j}
velocity at time 
v2 = 9.8\sqrt{3}\hat{i}  + (9.8 – 9.8*t)\hat{j}
v1 and v2 is perpendicular then
v1.v2 = 0
9.8*9.8*3 + 9.8*9.8 – 9.8*9.8t = 0
3 + 1 – t = 0
So t = 4sec
Hope it clears. If you like answer then pleas approve it.